1. Which electron transition in a hydrogen atom is associated with the highest frequency emission of energy?

a. n=3 to n=1

b. n=5 to n=4

c. n=4 to n=7

d. n=3 to n=2

AND

2. According to the following reaction, how many milliliters of 0.632 M FeCl3 are needed to completely react with 30.00 mL of 0.377 M Na2S ?

(molar mass of Fe2S3 = 207.88 g/mol)

3 Na2S(aq) + 2 FeCl3(aq) → Fe2S3(s) + 6 NaCl(aq)
Answer #1

1)
Emission is when electron moves from higher energy state to lower energy state.
This is true for a, b and d
Energy difference is more for a
So, frequency will be more for a
Answer: a

2)
Balanced chemical equation is:
3 Na2S + 2 FeCl3 ---> Fe2S3 + 6 NaCl


Here:
M(Na2S)=0.377 M
M(FeCl3)=0.632 M
V(Na2S)=30.0 mL

According to balanced reaction:
2*number of mol of Na2S =3*number of mol of FeCl3
2*M(Na2S)*V(Na2S) =3*M(FeCl3)*V(FeCl3)
2*0.377 M *30.0 mL = 3*0.632M *V(FeCl3)
V(FeCl3) = 11.93 mL

Answer: 11.93 mL